OpenStudy (anonymous):

CAN SOMEONE REALLY HELP ME WITH THESE SLOPE QUESTIONS 4. Determine the slope between the points (-1, 0) and (-2, 3)

OpenStudy (anonymous):

slope = rise over run, which is "change in y coordinates divided by change in x coordinates" using your two points as (x1, y1) and (x2, y2), slope = (y2 - y1) / (x2 - x1) You just put your actual coordinates in, subtract each part, then divide.

OpenStudy (anonymous):

so i just write this (3-0) / (-2-3) (0)/(1)

OpenStudy (anonymous):

@JakeV8 ummm so?

OpenStudy (anonymous):

I think you mixed up the part on the bottom... (3-0) was correct for the top. The bottom should have been -2 -(-1).

OpenStudy (anonymous):

ok so its -3?

OpenStudy (anonymous):

I get: (3-0) / (-2 - (-1)) = 3 / (-1) = -3 Is that how you did it too? If so, that's correct.

OpenStudy (anonymous):

yeah so wat u do after tht?

OpenStudy (anonymous):

@JakeV8 ?

OpenStudy (anonymous):

That's all there is... you found slope = -3. That's all this problem is asking you.

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is this the same thing too? 5.Determine the slope between the points (-1, -5) and (5, 4)

OpenStudy (anonymous):

yes, same thing... just repeat with those points. Be careful subtracting negatives though... that's where errors start. Give it a try and I'll check it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait i think this is wrong look at the numbers slope = rise over run (x1, y1) and (x2, y2), slope = (y2 - y1) / (x2 - x1) (3-0) / (-2 - (-1)) = 3 / (-1) = -3

OpenStudy (anonymous):

why does it seem wrong?

OpenStudy (anonymous):

because 4. Determine the slope between the points (-1, 0) and (-2, 3)

OpenStudy (anonymous):

x1=-1 y1=0 x2=-2 y2=3

OpenStudy (anonymous):

oops!!! sorry, my drawing is all wrong :(

OpenStudy (anonymous):

that's embarrassing... let me fix it.

OpenStudy (anonymous):

lol so was my answer wrong or right?

OpenStudy (anonymous):

I think I need a snack or something... my brain appears to have quit :(

OpenStudy (anonymous):

so is this right so far? (x1, y1) and (x2, y2), slope = (y2 - y1) / (x2 - x1) (4-5)/(5-(-1))

OpenStudy (anonymous):

(-1),(4)

OpenStudy (anonymous):

One more try on your #4, no drawings... slope between (-1, 0) and (-2, 3) is: rise / run = (3 - 0) / (-2 - (-1) ) = 3 / (-1) = -3 It goes up 3 for every one unit it moves to the left. That's correct... sorry for the drawing confusion.

OpenStudy (anonymous):

I think you messed up on the negatives... 5.Determine the slope between the points (-1, -5) and (5, 4) rise / run = (4 - (-5) ) / (5 - (-1) )

OpenStudy (anonymous):

(9),(6)?

OpenStudy (anonymous):

yes, but remember, it's a fraction, not a point... so you have slope = 9/6 You can reduce that fraction to 3/2. Slope is like a rule that says "go up 3 for every 2 you go to the right". Or, "rise 3 and run 2".

OpenStudy (anonymous):

If you "rise 3 and run 2" three times in a row, you can step from (-1,-5) up to (5,4).

OpenStudy (anonymous):

yeah so u just divide it by 3

OpenStudy (anonymous):

@JakeV8 so the slope = 3/2

OpenStudy (anonymous):

yes, that's correct.

OpenStudy (anonymous):

ok lol was the first question right too?

OpenStudy (anonymous):

yes, it was just my drawing that was messed up.

OpenStudy (anonymous):

ok what about this one Find the slope of the line graphed below.

OpenStudy (anonymous):

@JakeV8 did u get it?

OpenStudy (anonymous):

yes, I see it. What do you think it is? Rise over run... you just pick some point to start at and count the rise and run to some other point. An easy way is to choose where it crosses the x-axis on the left side, and count rise and run up to where it crosses the y-axis in the middle.

OpenStudy (anonymous):

(1,0)?

OpenStudy (anonymous):

Um, no... remember slope is a number, often a fraction... it isn't an ordered pair. If you want, you could calculate it like we did earlier... choose those same 2 points... (-3,0) and (0,5) then do rise over run.

OpenStudy (anonymous):

oh those are two points oon the line >.<

OpenStudy (anonymous):

@JakeV8 what you do after you find those points?

OpenStudy (anonymous):

you can find slope, just like you did earlier... slope = rise/run = (5-0) / (0 - (-3))

OpenStudy (anonymous):

3-0/-2+1=slope slope=-3

OpenStudy (anonymous):

I don't follow where you got those points. I think the easiest points to use are (-3,0) and (0,5), so slope = rise/run = (5-0) / (0 - (-3))

OpenStudy (anonymous):

We're talking about two or three different problems in this thread now... confusing... Maybe time for a new question posting if there are more slope questions.

OpenStudy (anonymous):

wait so it = (0,3)?

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

@pkjha3105 can u write out it out 4 me?

OpenStudy (anonymous):

slope=y2-y1/x2-x1 slope=3-0/-2+1 slope=-3

OpenStudy (anonymous):

@pkjha3105 ty ! can u help me with 2 more please :)